What is the angular resolution limit (degrees) set by diffraction for the 234-cm mirror diameter telescope (λ=560 nm)?
Answer #1

from Rayleigh criteria

x = 1.22 * L / D = 1.22* 560* 10^-9 / 2.34

x = 2.92 * 10^-7 rad

x = 1.674* 10^-5 deg

======